# Superconductivity Proximate to Non-Abelian Fractional Spin Hall Insulator in Twisted Bilayer MoTe2

### v1

@@ Discussion
In this manuscript, we theoretically uncover the remarkable emergence of superconductivity from strongly correlated topological flat bands, realized as an intermediate phase between the ferromagnetic Chern insulator and the non-Abelian fractional spin Hall insulator. This superconducting state is identified through multiple hallmarks, including negative binding energy, distinctive spectral features of the pair density matrix, and finite phase stiffness. Direct evaluation of the order parameter and the dominant wavefunction of the pair density matrix reveals its pairing symmetry to be time-reversal symmetric, extended nodal $s$-wave. The nodal character is further corroborated by the observation that the superfluid stiffness $D_{s}$ is of the same order as the binding energy $|E_{b}|$, reaching a maximum ratio of $D_{s}/\left(|E_{b}|/2\right)\approx 0.9$ in our calculations, in contrast to conventional gapped superconductors where $D_{s}\gg|E_{b}|/2$ [61, 62].
The experimental observation of both ferromagnetism [52] and the fractional spin Hall insulator [13, 14] in twisted bilayer MoTe2 strongly suggests that tuning between these regimes could open a pathway to realizing the intermediate superconducting phase uncovered in our analysis. More specifically, the effective screening length $d$ could be tuned through gating or substrate engineering, thereby promoting superconductivity. This follows from the two following observations. First, the length scale $d$ arises from band mixing and is roughly set by the inverse of the effective band gap between the first and second moiré bands [40]. Second, in twisted bilayer MoTe2, the Coulomb energy is comparable to the bare band gap, so at filling $\nu_{h}=3$ the effective band gap is strongly renormalized by interactions. As a result, by controlling interactions—and thereby the effective band gap—the parameter $d$ can be effectively controlled. Indeed, a rough estimate of $d$ for experimentally available devices [63] places it near the transition points [50], suggesting that superconductivity may be within experimental reach. On the theoretical side, it would be valuable to further clarify the role of non-uniform quantum geometry in stabilizing superconductivity [28, 29, 30, 31, 32, 33, 34, 35], as well as the relation between the intermediate superconductivity found here and anyon superconductivity [28, 29, 30, 31, 32, 33, 46, 47].
@@ Methods
We make use of spin $\mathrm{U}_{s}(1)$, moiré translation to divide the many-body Fock space into subspaces of fixed hole number, magnetization, and momentum. When applicable, we also apply $C_{2}^{y}$ and time-reversal symmetries. Within each subspace, the many-body Hamiltonian isrepresented with a sparse matrix, and diagonalized using Krylov methods. Using the exact diagonalization results, we also compute the intervalley entanglement for various many-body states. The intervalley entanglement entropy [Fig. 3(a)] is defined as
[…]
where $\hat{\rho}_{0,\uparrow}$ is the $K$-valley reduced density operator of the many-body ground state, obtained by partially tracing out the $K^{\prime}$-valley Hilbert space. Other quantities have been defined in the main text.
- Note Added: While finalizing this manuscript, we became aware of a similar result [64]. In this paper, the authors tuned the intervalley interaction in a different form from Eq. (2), and also found a superconductor between the ferromagnetic Chern insulator and fractional quantum spin Hall insulator at half-filling.
@@ Acknowledgments
We thank Fiona Burnell, Debanjan Chowdhury, Yong Baek Kim, Yves Kwan, Inti Sodemann, Ashvin Vishwanath, and Ya-Hui Zhang for helpful discussions. This work is financially supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2401-03, the NRF of Korea (Grants No. RS-2023-00208291, RS-2024-00410027, 2023M3K5A1094810, RS-2023-NR119931, RS-2024-00444725, RS-2023-00256050, IRS-2025-25453111) funded by the Korean Government (MSIT), the Air Force Office of Scientific Research under Award No. FA23862514026, and Institute of Basic Science under project code IBS-R014-D1. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. The work from DGIST was supported by the National Research Foundation of Korea (NRF) (Grant No. RS-2025-00557717, RS-2023-00269616) and the Nano and Material Technology Development Program through the National Research Foundation of Korea (NRF) funded by Ministry of Science and ICT (No. RS-2024-00444725). We also acknowledge the partner group program of the Max Planck Society.

### v2

@@ Discussion
In this manuscript, we theoretically uncover the remarkable emergence of superconductivity from strongly correlated topological flat bands, realized as an intermediate phase between the ferromagnetic Chern insulator and the non-Abelian fractional spin Hall insulator. This superconducting state is identified through multiple hallmarks, including negative binding energy, distinctive spectral features of the pair density matrix, and finite phase stiffness. Direct evaluation of the order parameter and the dominant wavefunction of the pair density matrix reveals its pairing symmetry to be time-reversal symmetric, extended nodal $s$-wave. We further elucidate two complementary pairing mechanisms: a Kohn–Luttinger instability arising from non-uniform quantum geometry when approached from the normal metallic side [24, 95, 74], and the condensation of charge-$e/2$ anyons when approached from the fractional spin Hall insulator. In particular, the latter mechanism establishes our system as a realistic platform for realizing anyon superconductivity [79, 80, 65, 86, 103, 59, 78, 30, 101, 44, 77, 73, 48, 49].
The experimental observation of both ferromagnetism [50] and the fractional spin Hall insulator [39, 38] in twisted bilayer MoTe2 strongly suggests that tuning between these regimes could open a pathway to realizing the intermediate superconducting phase uncovered in our analysis. More specifically, the effective screening length $d$ could be tuned through gating or substrate engineering, thereby promoting superconductivity. This follows from the two following observations. First, the length scale $d$ arises from band mixing and is roughly set by the inverse of the effective band gap between the first and second moiré bands [2]. Second, in twisted bilayer MoTe2, the Coulomb energy is comparable to the bare band gap, so at filling $\nu_{h}=3$ the effective band gap is strongly renormalized by interactions. As a result, by controlling interactions—and thereby the effective band gap—the parameter $d$ can be effectively controlled. Indeed, a rough estimate of $d$ for experimentally available devices [102] places it near the transition points [1], suggesting that superconductivity may be within experimental reach. Additionally, larger-scale numerical studies—such as quasi-one-dimensional density-matrix renormalization group calculations—would be valuable for further mitigating the finite-size effects inherent in our exact diagonalization results and for performing systematic scaling analyses near the anyon-condensation transition [1], thereby yielding deeper insight into the physics of this model.
@@ Methods
We make use of spin $\mathrm{U}_{s}(1)$, moiré translation to divide the many-body Fock space into subspaces of fixed hole number, magnetization, and momentum. When applicable, we also apply $C_{2}^{y}$, valley-reduced inversion [37, 3], and time-reversal symmetries. Within each subspace, the many-body Hamiltonian is represented with a sparse matrix, and diagonalized using Krylov methods.
We note that we estimate the superconducting coherence length as $\ell_{\text{SC}}\approx 1.83\,a_{M}$ [1], which is less than the side lengths of the finite lattices used in our calculations. As such, we expect superconducting pairing to be visible, despite finite-size effects [1].
- Note Added: While finalizing this manuscript, we became aware of a paper with some overlapping results [18]. In this paper, the authors tuned the intervalley interaction in a different form from Eq. (2), and also found a superconductor between the ferromagnetic Chern insulator and non-Abelian fractional spin Hall insulator at half-filling.
@@ Acknowledgments
We thank Fiona Burnell, Debanjan Chowdhury, Hee-Cheol Kim, Yong Baek Kim, Yves Kwan, Inti Sodemann, Ashvin Vishwanath, and Ya-Hui Zhang for helpful discussions. This work is financially supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2401-03, the NRF of Korea (Grants No. RS-2024-00410027, RS-2023-NR119931, RS-2024-00444725, RS-2023-00256050, IRS-2025-25453111, RS-2025-08542968) funded by the Korean Government (MSIT), the Air Force Office of Scientific Research under Award No. FA23862514026, and Institute of Basic Science under project code IBS-R014-D1. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. The work from DGIST was supported by the National Research Foundation of Korea (NRF) (Grant No. RS-2025-00557717, RS-2023-00269616) and the Nano and Material Technology Development Program through the National Research Foundation of Korea (NRF) funded by Ministry of Science and ICT (No. RS-2024-00444725). We also acknowledge the partner group program of the Max Planck Society. Part of this work was supported by Global Partnership Program of Leading Universities in Quantum Science and Technology (RS-2025-02317602)
@@
Author Contributions
G. Y. C. conceived and supervised the project. C. A. performed the Hartree–Fock analysis, exact diagonalization calculations on moiré bands, and numerical characterization of the many-body ground-states. D. S. carried out the effective field-theory analysis and the analysis of anyon data related to anyon condensation. G. L. performed exact diagonalization calculations on generalized first Landau levels. Y. K. provided experimental inputs that informed the theoretical proposal. C. A., D. S., and G. Y. C. co-wrote the manuscript. All authors discussed the results and contributed to the final version of the manuscript.